### v1

# TBStar-Edit: From Image Editing Pattern Shifting to Consistency Enhancement

Abstract
Recent advances in image generation and editing technologies have enabled state-of-the-art models to achieve impressive results in general domains. However, when applied to e-commerce scenarios, these general models often encounter consistency limitations. To address this challenge, we introduce TBStar-Edit, an new image editing model tailored for the e-commerce domain. Through rigorous data engineering, model architecture design and training strategy, TBStar-Edit achieves precise and high-fidelity image editing while maintaining the integrity of product appearance and layout. Specifically, for data engineering, we establish a comprehensive data construction pipeline, encompassing data collection, construction, filtering, and augmentation, to acquire high-quality, instruction-following, and strongly consistent editing data to support model training. For model architecture design, we design a hierarchical model framework consisting of a base model, pattern shifting modules, and consistency enhancement modules. For model training, we adopt a two-stage training strategy to enhance the consistency preservation: first stage for editing pattern shifting, and second stage for consistency enhancement. Each stage involves training different modules with separate datasets. Finally, we conduct extensive evaluations of TBStar-Edit on a self-proposed e-commerce benchmark, and the results demonstrate that TBStar-Edit outperforms existing general-domain editing models in both objective metrics (VIE Score) and subjective user preference.
Figure 1: Overview of TBStar-Edit. TBStar-Edit can handle a variety of tasks in both general and e-commerce domains.
1 Introduction
Driven by the rapid advancements in Diffusion Models Ho et al. (2020); Lipman et al. (2022) and text-to-image generation Esser et al. (2024); bla (2024), AI-powered image editing technologies have experienced significant progress in recent years. From early UNet-based models such as Instruct-pix2pix Brooks et al. (2023), EmuEdit Sheynin et al. (2024), and HIVE Zhang et al. (2024) to more recent DiT-based models Peebles and Xie (2023) like ACE Han et al. (2024), ACE++ Mao et al. (2025), and Step1X-Edit Liu et al. (2025), there have been substantial breakthroughs in terms of model parameters, the quantity and quality of training data, and the resolution of generated images. Currently, editing models feature increasingly versatile editing capabilities and noticeably improved editing performance. Numerous companies have introduced multimodal editing models, including GPT-4o Hurst et al. (2024), Nano Banana Comanici et al. (2025), Flux-Kontext Batifol et al. (2025), Qwen-Image-Edit Wu et al. (2025), and SeedEdit Wang et al. (2025); Seedream et al. (2025). These general-purpose editing models exhibit powerful abilities in complex semantic understanding, multilingual scenarios, and creative editing, making rapid image modification a practical reality.
Although these models have demonstrated impressive performance in general domains, their performance still limited in the e-commerce scenarios. The primary reason for this limitation is inadequate consistency preservation, which often manifests as unintended changes to the main subject or background when local modifications are made. Consequently, discrepancies may arise in the identity of the product or in fine-grained details between the edited and original images. To address this issue, we propose TBStar-Edit, a model specifically designed for e-commerce scenarios requiring high consistency. By employing rigorous data engineering, model architecture, and training strategies, TBStar-Edit enables precise and high-fidelity image editing while maintaining the integrity of the product and its layout.
For data engineering, we observed that existing open-source datasets such as HQEdit Hui et al. (2024), UltraEdit Zhao et al. (2024), OmniEdit Wei et al. (2024), and GPT-Image-Edit Wang et al. (2025) suffer from limitations including imprecise instruction-following, insufficient consistency preservation, and suboptimal image quality. Relying solely on these open-source datasets makes it challenging for models to learn high-consistency, instruction-driven editing effectively. To address these shortcomings, we propose a comprehensive data construction pipeline consisting of four distinct phases: data collection, editing pairs construction, data filtering, and post-processing. Each phase is meticulously designed: realism data is prioritized in the collection phase; four task-specific methods are introduced in the editing pairs construction phase; a combination of vision-language models and manual quality inspection is employed during data filtering; and editing instructions are diversified in the data augmentation phase. Through this pipeline, precise instruction-following, strong subject consistency, and high image quality are ensured in the constructed dataset to support consistency enhancement.
For model architecture, we design a hierarchical framework consisting of a base model, pattern shifting modules, and consistency enhancement modules. This layered structure enables our model to flexibly adapt to various backbone models and aligns with our training strategy. For training strategy, we divide the process into two stages: first acquisition of editing capability and second enhancement of consistency. In each stage, distinct datasets are used to separately train the pattern shifting module and the consistency enhancement module. This approach first transforms a T2I generation model into a TI2I editing model, and subsequently reinforces consistency within the shifted editing model.
Finally, we conducted a comprehensive evaluation on our self-constructed e-commerce benchmark, Ecom-Edit Bench, comparing TBStar-Edit with state-of-the-art general-purpose models including Flux-Kontext, Qwen-Image-Edit, NanoBanana, and Seedream 4.0 Seedream et al. (2025). The evaluation covered instruction-following, consistency preservation, and image quality. The visualization results, VIE Score Ku et al. (2023) calculations, and user preference comparisons demonstrate that TBStar-Edit outperform existing general-domain state-of-the-art models in both objective metrics and subjective assessments in high-consistency e-commerce scenarios, validating TBStar-Edit’s strong capability in consistency preservation.
Figure 2: Pipeline of data engineering.
2 Data Engineering
The performance of image editing models is highly dependent on the training data. To enable a model to learn precise and high-fidelity image modifications, the training data must possess three key characteristics: accurate instruction following, strong subject consistency, and high image quality. However, current open-source datasets suffer from significant issues such as severe detail distortion and identity (ID) drift. To address these problems, we have constructed a complete data construction pipeline. This pipeline consists of four stages: data collection, editing pair construction, data filtering, and data post-processing. The overall datapipeline is illustrated in Figure 2.
2.1 Data Collection
Prior work has often incorporated a large volume of synthetic data during data collection. While this approach significantly reduces the difficulty of data acquisition and enhances data diversity, an over-reliance on synthetic data can diminish the realism of editing results, thereby compromising the model’s performance in real-world applications. Therefore, we have minimized the proportion of generated images in our source data, opting instead to collect a large number of real-world images. To ensure the diversity of our source data, we gathered images from a variety of scenes and sources, including general-domain portraits and object images from the web and open-source datasets, as well as internal e-commerce portraits and product images. Lastly, to guarantee the quality of the source images, we applied a filtering process based on resolution and Aesthetic Scores.
2.2 Editing Pairs Construction Methods
Constructing editing pairs from collected raw image data is the core process in building image editing datasets. Given the wide variety of image editing tasks, each with distinct characteristics and requirements, we systematically categorize these tasks and, accordingly, establish four different methods for constructing editing pairs: the Editing Expert Model Based Method, the Template Based Method, the In-Context Generation Based Method, and the LoRA Based Method.
Figure 3: Details of Auxiliary Information and In-context Generation in editing pairs construction.
2.2.1 Editing Expert Model Based Method
Although numerous models have demonstrated exceptional performance on various editing tasks, their applicability is often restricted to specific domains, thus limiting their generalizability. To harness the specialized strengths of these models and integrate their capabilities, we introduce a novel pipeline that employs a Vision-Language Model (VLM) for instruction generation and Task-specific Editing Experts for image manipulation.
In the context of instruction generation, previous approaches typically involve directly feeding images into a VLM. This method, however, is susceptible to the VLM’s inherent biases, resulting in generated data with poor generalization. For instance, due to a bias towards object removal, a VLM may predominantly generate instructions for excising objects located at the image periphery. To mitigate such biases, we augment the VLM’s input with auxiliary information. This information, provided alongside the original image, is twofold: (1) a list of object categories present in the image and their respective counts, and (2) a set of pre-defined, task-relevant instructions. The specific format of this auxiliary information is illustrated in Figure 3(a). By conditioning the VLM on this data, we aim to reduce its biases and enhance the diversity and generalizability of the generated instructions.
Subsequently, the original image and the VLM-generated instruction are processed by a Task-specific Editing Expert—which may comprise open-source models, proprietary workflows, or other pipelines—to produce the final edited image. In the final stage, both the original and edited images are fed back into the VLM to generate more fine-grained and comprehensive instructions. This methodology, through the strategic incorporation of auxiliary information and the use of specialized editing experts, enables the efficient generation of a large-scale, highly generalizable dataset, thereby facilitating rapid performance gains on targeted editing tasks.
2.2.2 Template Based Method
For certain editing tasks that can be templated, such as watermark removal, text element removal, text modification, and lighting/shadow adjustment, we employ a template-based data construction approach. The main workflow of this method involves: (1) generating edited images by adding image templates to the original image, applying 3D rendering templates, or modifying specific template elements in relevant regions of the original image; (2) generating editing instructions either through predefined rules or manual input, followed by enhancement of instruction diversity using a VLM. As this approach ensures that regions outside the templates remain completely unchanged, the resulting data exhibits excellent subject consistency
2.2.3 In-Context Generation Based Method
Although the Editing Expert-based and Template-based data construction methods address most editing tasks, not all tasks have corresponding editing experts or can be templated. For tasks that lack editing experts and cannot be templated—such as changes in a person’s age or physique—we adopt the In-Context Generation Based Method. This approach fully leverages the in-context generation capability of T2I models and, through effective template design, enables the model to generate image pairs or sequences with strong identity preservation, as illustrated in Figure 3(b). Subsequently, these image pairs or sequences are split, and editing instructions are generated using a Vision-Language Model (VLM). This method has a broad range of applications and is especially well-suited as a supplementary data construction strategy for specialized tasks.
2.2.4 LoRA Hu et al. (2022) Based Method
For specialized editing tasks that lack editing experts and cannot be templated, we employ an alternative approach by utilizing LoRA for less-to-more data generation. Initially, a small number of editing pairs (ranging from several dozen to a few hundred) are created for these specific tasks, either through web scraping or manual annotation. These pairs are then used to train a task-specific LoRA model. Once LoRA training is complete, we input a large set of raw image-instruction pairs—annotated by VLM—into the task-specific LoRA model, which infers and produces a substantial volume of edited images. Subsequently, both the original and edited images are input into the VLM for refined instruction generation. Compared to the In-Context Generation Based Method, this approach better ensures the authenticity of the data and helps improve inference performance in real-world scenarios, although it demands more time and human resources.
2.3 Data Filter and Data Postprocess
Although our editing data construction methods are rigorously designed to ensure high quality and a high acceptance rate for edited images, limitations inherent to various editing models and persistent biases in VLMs cannot be entirely eliminated. As a result, a certain proportion of low-quality edited data inevitably remains in our constructed dataset, potentially affecting model training performance. To further guarantee the quality of the editing data, we implement a strict filtering process for the constructed editing pairs. First, all constructed data are evaluated by a VLM, which scores them across multiple dimensions, including instruction-following, consistency between the original and edited images, and image quality. Low-scoring data are excluded based on these assessments. Furthermore, due to discrepancies between VLM and human judgment regarding the quality of edited data, we conduct manual quality inspection to obtain a set of high-quality data.
Finally, To ensure precise model responses to diverse instructions in both Chinese and English, we enhanced the editing instructions in two ways. First, original instructions were translated into Chinese. Then, a large language model was used to generate various forms—including synonyms, interrogatives, and passive constructions—for both languages. All original and augmented instructions were collected for flexible use in model training.
Figure 4: Framework of the TBStar-Edit model.
3 TBStar-Edit Model
TBStar-Edit is built upon two key principles: (1) fully leveraging the powerful in-context generation capability of T2I model and transforming it into image editing capability; (2) focusing the capabilities of the converted editing model to enhance subject consistency in edited images, ensuring practical deployment in real-world scenarios. To cover these principles, we design a flexible, hierarchical model framework that can be adapted to various T2I backbone models, along with a training strategy that first transforms a T2I generation model into a TI2I editing model, and subsequently enhance consistency within the shifted editing model.
3.1 Model Architecture
Our model adopts a hierarchical framework comprising a base model, pattern shifting modules, and consistency enhancement modules, enabling high flexibility and scalability. This design facilitates seamless adaptation to various open-source or proprietary backbone models, such as Flux, Kontext, HiDream, and Qwen-Image. For each DiT-Block, a LoRA layer is inserted at the QKV Attention position as the pattern shifting module, and another two LoRA layers are added at the Projections following Attention as the consistency enhancement module. During training, all original parameters of the backbone model are frozen, and only the parameters of the pattern shifting and consistency enhancement modules are optimized.
We implemented editing adaptation and capability focusing based on Flux, Kontext, HiDream, and Qwen-Image. Notably, Qwen-Image demonstrates excellent comprehension and generation in Chinese contexts, especially for Chinese e-commerce text and complex instruction parsing. Consequently, we selected Qwen-Image as the base model for TBStar-Edit.
3.2 Training Strategy
Multi-stage Training Strategy.
The primary objectives of our training are: (1) to instill foundational editing paradigms within the Text-to-Image (T2I) generative model, (2) to refine the model’s capabilities, focusing on enhancing subject consistency. Consequently, our training is structured into two distinct phases: Stage 1 for Pattern Shifting and Stage 2 for Consistency Enhancement. Stage 1: Pattern Shifting. To equip the model with broad-spectrum editing capabilities, we train it on a composite dataset comprising large-scale open-source data and our high-quality constructed data. During this phase, all LoRA parameters are made trainable, including both the pattern shifting modules and consistency enhancement modules. Stage 2: Consistency Enhancement. This stage is designed to specialize the model for important editing tasks in the e-commerce domain and aim to address the consistency issues left from stage 1. Crucially, to ensure the general editing abilities acquired in the previous stage are largely retained, we train solely on high-quality proprietary datasets, with the pattern shifting modules freezed and only the consistency enhancement modules updated.
Mixed Chinese-English Instruction Training.
To improve the model’s precise response to instructions in different languages and forms, as well as to enhance its editing performance under diverse instructions—especially for Chinese instructions and the consistency of Chinese editing—the model samples both original and augmented Chinese and English instructions from the training dataset at a certain ratio. Training is then conducted using this mixed set of instructions.
Figure 5: visualization comparison between state-of-the-art models in general domain.
4 Experiment
4.1 Visual Comparison
The visualization results for several tasks are shown in Figure 5. It can be observed that TBStar-Edit achieves strong performance not only on general tasks such as object replacement and object removal, but also on e-commerce-specific scenarios including portrait background replacement and lighting adjustment. The edited images demonstrate a high degree of balance between instruction following and subject consistency. Compared with other models, TBStar-Edit excels particularly in consistency; while maintaining faithful adherence to the input instructions, it delivers best-in-class results on identity preservation, action retention, and detail maintenance in portrait editing among all evaluated models.
4.2 Quantitative Evaluation
To assess the model’s performance on e-commerce editing tasks, we established an e-commerce evaluation benchmark, EcomEdit-Bench, inspired by widely adopted community standards such as GEdit-Bench. This benchmark encompasses a variety of common e-commerce scenarios, including portrait pose editing, background replacement, accessory addition, hairstyle modification, addition, removal, and replacement of e-commerce elements, text deletion or modification, watermark removal, lighting adjustment, and more, with a total of 380 image-instruction pairs. We employ an automated and objective evaluation protocol to measure the model’s effectiveness. For each generated image, GPT-4o is used to score three aspects: instruction adherence, image quality, and consistency and calculate the average VIE Score .
We evaluated the e-commerce editing performance of several mainstream editing models on the EcomEdit-Bench benchmark. Qwen-Image-Edit-2509 and Flux-Kontext were evaluated via local deployment using open-source code, while NanoBanana and Seedream4.0 were assessed via API calls (with some test cases excluded due to security checks). GPT-Image-1 was not included in the quantitative evaluation, as it is only capable of generating images in three different resolutions and exhibits relatively poor consistency. The objective evaluation results, measured by the VIE Score, are presented in Table . As shown, TBStar-Edit achieves the highest overall and Semantic Consistency scores among all compared models. These results demonstrate that TBStar-Edit effectively balances instruction adherence and consistency in editing tasks, making it particularly suitable for e-commerce and other application domains that demand high consistency.
Table 1: Evaluation results on EcomEdit-Bench. $G_{SC}$: Semantic Consistency, $G_{PQ}$: Perceptual Quality, $G_{O}$: Overall. Higher is better. Bold font indicates the highest value.
Model $G_{SC}$ $G_{PQ}$ $G_{O}$
Flux-Kontext-Dev 7.552 8.441 7.732
HiDream-E1 6.849 7.701 6.658
SeedEditv3.0-en 8.701 8.567 8.376
SeedEditv3.0-cn 8.757 8.595 8.410
NanoBanana 8.217 8.775 8.087
Qwen-Image-Edit-2509-en 8.971 8.617 8.579
Qwen-Image-Edit-2509-cn 8.954 8.630 8.569
Seedream4.0-en 9.002 8.701 8.643
Seedream4.0-cn 9.041 8.756 8.714
TBStar-Edit-en 9.139 8.575 8.688
TBStar-Edit-cn 9.206 8.579 8.746
Table 2: Human preference evaluation.
Model TBStar-Edit win TBStar-Edit loss Tie
TBStar-Edit vs Flux-Kontext-Dev 55.80% 18.90% 25.30%
TBStar-Edit vs Qwen-Image-Edit-2509 46.94% 26.53% 28.57%
TBStar-Edit vs NanoBanana 44.20% 32.50% 23.40%
TBStar-Edit vs Seedream4.0 40.00% 28.89% 31.11%
Although the VIE Score provides a comprehensive evaluation of instruction adherence, consistency preservation, and image quality, it is not fully accurate due to two inherent limitations of VLMs: (1) VLMs struggle to identify and compare subtle features, often resulting in less precise consistency assessment; (2) VLMs are prone to hallucinations, occasionally leading to imprecise judgments regarding instruction following. Therefore, relying solely on the VIE Score to evaluate editing model performance is insufficient. To address this, we additionally conducted a human evaluation on the editing results produced by our model on EcomEdit-Bench. In the evaluation process, subjects were presented with the original image, the corresponding English and Chinese instructions, and two edited images—one generated by TBStar-Edit and the other randomly selected from Flux-Kontext, Qwen-Image-Edit-2509, Seedream4.0, or NanoBanana. The presentation order was randomized. Participants then selected the better edited image according to three criteria: instruction adherence, image quality, and consistency preservation. The results of the human evaluation are shown in Table . Compared with state-of-the-art models in the general domain, such as NanoBanana, Seedream4.0, and Qwen-Image-Edit-2509, TBStar-Edit achieved a higher preference rate, indicating its competitive advantage in e-commerce scenarios that demand high consistency.
5 Conclusion
In this report, a new model named TBStar-Edit was proposed, focusing on high product-level consistency for e-commerce applications. Accurate and faithful image editing was achieved, resulting in significant improvements in model performance within high-consistency e-commerce scenarios. To address the issue of low-quality open-source data, a comprehensive data construction pipeline was introduced, encompassing data collection, construction, filtering, and augmentation. Through this pipeline, a large amount of high-quality data was generated, characterized by precise instruction following, strong object consistency, and superior image quality. Based on the concept of enhancing editing capability before reinforcing consistency, a hierarchical model framework was developed, consisting of a base model, pattern shifting modules, and consistency enhancement modules. A two-stage training strategy was adopted, enabling T2I models to acquire editing abilities and substantially improve consistency. Experimental results demonstrated that, in high-consistency e-commerce scenarios, the proposed model significantly outperformed current general SOTA models.